# Directional Convergence, Benign Overfitting of Gradient Descent in leaky ReLU two-layer Neural Networks

### v2

@@ Abstract
In this paper, we study benign overfitting of fixed width leaky ReLU two-layer neural network classifiers trained on mixture data via gradient descent. We provide both, upper and lower classification error bounds, and discover a phase transition in the bound as a function of signal strength. The lower bound leads to a characterization of cases when benign overfitting provably fails even if directional convergence occurs. Our analysis allows us to considerably relax the distributional assumptions that are made in existing work on benign overfitting of leaky ReLU two-layer neural network classifiers. We can allow for non-sub-Gaussian data and do not require near orthogonality. Our results are derived by establishing directional convergence of the network parameters and studying classification error bounds for the convergent direction. Previously, directional convergence in (leaky) ReLU neural networks was established only for gradient flow. By first establishing directional convergence, we are able to study benign overfitting of fixed width leaky ReLU two-layer neural network classifiers in a much wider range of scenarios than was done before.
@@ 1 Introduction
Our results are derived by establishing directional convergence of the network parameters with precise characterization of the convergent direction and by obtaining a classification error bound of that direction. Due to the difficulty in establishing directional convergence of gradient descent for ReLU type networks, prior work for leaky ReLU networks were limited to either gradient flow (Frei et al. (2023a)), smoothed approximation of leaky ReLU (Frei et al. (2022)), or nearly orthogonal data regime (Xu and Gu (2023)), i.e. the predictors $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,n}\subset\mathbb{R}^{p}$ satisfy $\|\boldsymbol{x}_{i}\|^{2}\gg\max_{i\neq k}|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle|$.
@@ 2 Related Work
More closely related work is Frei et al. (2022), which studied benign overfitting in smoothed leaky ReLU two-layer neural networks on mixture data. Their work was extended by Xu and Gu (2023) to non-smooth leaky ReLU like activation. While Xu and Gu (2023)’s work is not limited to (leaky) ReLU, their work is limited to the nearly orthogonal data regime and only holds for isotropic mixture with bounded log-Sobolev constant, which is a stronger assumption than sub-Gaussian mixture. Our work establishes benign overfitting beyond the nearly orthogonal data regime and can be applied to a wider class of mixtures than the prior work such as polynomially tailed mixture. Moreover, we also establish a novel lower bound of the classification error in the case of Gaussian mixture, which shows tightness of our result and helps characterize cases when benign overfitting provably fails. Even with gradient flow, benign overfitting in leaky ReLU neural networks is limited to nearly orthogonal data setting and sub-Gaussian mixture model (Frei et al. (2023a)).
@@
Most prior work on the implicit bias of (leaky) ReLU neural networks for binary classification has considered gradient flow. For gradient flow, Lyu and Li (2020) showed that any limit point of the network parameter of homogeneous neural networks trained with exponential type loss is a KKT point of a constrained optimization problem and that margin maximization occurs. Ji and Telgarsky (2020) further established directional convergence of the parameter of homogeneous neural networks trained with the exponential type loss. Taking advantage of the directional convergence of gradient flow, Lyu et al. (2021); Bui Thi Mai and Lampert (2021); Sarussi et al. (2021); Frei et al. (2023b); Min et al. (2024) obtained detailed characterizations of the convergent direction of the parameters of (leaky) ReLU two-layer neural networks under further assumptions on the training data.
In contrast to gradient flow, directional convergence of network parameters for ReLU type neural networks was not proven for gradient descent with exponential type loss, even under strong assumptions. As a result, only weaker forms of implicit bias have been obtained for gradient descent. Lyu and Li (2020) showed that margin maximization also occurs for the network parameters of a homogeneous network trained by gradient descent, but their assumptions include smoothness of the network, and hence rule out (leaky) ReLU networks. Frei et al. (2023b) showed that the stable rank of network parameters stays bounded by a constant for two-layer neural networks, but their result is limited to the smoothed approximation of leaky ReLU. Kou et al. (2023b) proved similar results for (leaky) ReLU two-layer neural networks assuming that the number of positive and negative neurons are equal. Moreover, Frei et al. (2023b); Kou et al. (2023b) are limited to the rather specialized scenario of nearly orthogonal data regime. Cai et al. (2025) showed directional convergence of gradient descent, but their result is not applicable to ReLU type neural networks since their result requires networks to be twice differentiable with respect to the parameters. Schechtman and Schreuder (2025) extended Lyu and Li (2020) by removing twice differentiability but their result does not guarantee directional convergence and assumes that training is already in the late-stage.
@@
Gradient Flow Ji and Telgarsky (2020) Frei et al. (2023a) NA
@@ 3 Preliminaries
where $\boldsymbol{x}\in\mathbb{R}^{p}$, $\boldsymbol{w}_{j}\in\mathbb{R}^{p}$, $W=(\boldsymbol{w}_{1},\boldsymbol{w}_{2},\ldots,\boldsymbol{w}_{m})\in\mathbb{R}^{p\times m}$, $|a_{j}|=\pm\tfrac{1}{\sqrt{m}}$ and $\phi$ is $\gamma$-leaky ReLU activation, i.e. $\phi(x)=\max\{x,\gamma x\}$ for $\gamma\in(0,1)$. Letting $\zeta(x)=1$ if $x\geq 0$ and $\zeta(x)=\gamma$ if $x<0$, we may conveniently write $\phi(x)=\zeta(x)x$.
@@ 4 Directional Convergence of Gradient Descent
We start from our main result on directional convergence of gradient descent in leaky ReLU two-layer neural networks trained on mixture data. Our result provides precise characterization of the convergent direction, which is a key for our analysis on benign overfitting.
@@
Assumption 4.4 and 4.6 allow us to carefully control $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle$. Small $\varepsilon_{1},\varepsilon_{3}$ ensure that the training data are nearly orthogonal. By taking $\boldsymbol{\mu}=0$, this case includes the nearly orthogonal data regime studied in the prior work (Frei et al. (2023b); Kou et al. (2023b)). The other assumptions are analogous to Case 1.
@@
Furthermore, the convergent direction $\{\boldsymbol{\hat{w}}_{j}\}_{j\in J}$ can also be given by $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{+}$ for $j\in J_{+}$ and $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{-}$ for $j\in J_{-}$, which are the solution to the following optimization problem:
@@
We note that condition (i) goes beyond nearly orthogonal data regime studied in the existing literature since there is no upper bound to the magnitude of the signal $\boldsymbol{\mu}$, while condition (ii) contains the nearly orthogonal data regime studied in Frei et al. (2023b); Kou et al. (2023b). As discussed in Introduction, directional convergence of gradient descent in ReLU type neural networks was not established in the existing literature even under nearly orthogonal data regime. In addition, the characterizations of the convergent direction in Theorem 4.8 were obtained only for gradient flow trained on nearly orthogonal data by Frei et al. (2023b).
@@ 6 Benign Overfitting on Mixture Models
    $n\geq C$, $\alpha\leq\frac{1}{8{\rm tr}(\Sigma)}$, $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$, $\|\boldsymbol{\mu}\|^{2}\geq C\|\Sigma^{]}frac{1}{2}\boldsymbol{\mu}\|$,
@@
Theorem 6.2 implies that benign overfitting occurs if $n\|\boldsymbol{\mu}\|^{4}\gg\|\Sigma\|{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\gg\|\Sigma\|$ under (b) is satisfied. For Gaussian mixture model, we see that benign overfitting provably fails even though directional convergence occurs if $n\|\boldsymbol{\mu}\|^{4}\lesssim\beta_{min}{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\lesssim\beta_{min}$ under (b).
@@
Consider model (PM) with $\Sigma=I_{p}$. Then, for sufficiently large $n$, benign overfitting occurs with the convergent direction $\hat{W}$ with high probability if either one of the following holds:
@@ 7 Proof Sketch
The key to show directional convergence of the gradient descent is activation of neurons. We say that $j$-th neuron $\phi(\langle\cdot,\boldsymbol{w}_{j}\rangle)$ is activated if $a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle)>0$ holds for any $i\in I$, which can be equivalently written as $a_{j}y_{i}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle>0$ for any $i\in I$.
@@ Appendix B Proof of Directional Convergence
$=\mathbb{P}(|\langle\boldsymbol{w},\boldsymbol{z}\rangle|>\langle\boldsymbol{w},\boldsymbol{\mu}\rangle).$

### v3

@@ Abstract
In this paper, we provide sufficient conditions of benign overfitting of fixed width leaky ReLU two-layer neural network classifiers trained on mixture data via gradient descent. Our results are derived by establishing directional convergence of the network parameters and classification error bound of the convergent direction. Our classification error bound also lead to the discovery of a newly identified phase transition. Previously, directional convergence in (leaky) ReLU neural networks was established only for gradient flow. Due to the lack of directional convergence, previous results on benign overfitting were limited to those trained on nearly orthogonal data. All of our results hold on mixture data, which is a broader data setting than the nearly orthogonal data setting in prior work. We demonstrate our findings by showing that benign overfitting occurs with high probability in a much wider range of scenarios than previously known. Our results also allow us to characterize cases when benign overfitting provably fails even if directional convergence occurs. Our work thus provides a more complete picture of benign overfitting in leaky ReLU two-layer neural networks.
@@ 1 Introduction
Our results are derived by establishing directional convergence of the network parameters with precise characterization of the convergent direction and by obtaining a classification error bound of that direction. Due to the difficulty in establishing directional convergence of gradient descent for ReLU type networks, prior work for leaky ReLU networks were limited to either gradient flow (Frei et al. (2023b)), smoothed approximation of leaky ReLU (Frei et al. (2022)), or nearly orthogonal data regime (Xu and Gu (2023)), i.e. the predictors $\{\boldsymbol{x}_{i}\}_{i=1,\ldots,n}\subset\mathbb{R}^{p}$ satisfy $\|\boldsymbol{x}_{i}\|^{2}\gg\max_{i\neq k}|\langle\boldsymbol{x}_{i},\boldsymbol{x}_{k}\rangle|$.
@@ 2 Related Work
More closely related work is Frei et al. (2022), which studied benign overfitting in smoothed leaky ReLU two-layer neural networks on mixture data. Their work was extended by Xu and Gu (2023) to non-smooth leaky ReLU like activation. While Xu and Gu (2023)’s work is not limited to (leaky) ReLU, their work is limited to the nearly orthogonal data regime and only holds for isotropic mixture with bounded log-Sobolev constant, which is a stronger assumption than sub-Gaussian mixture. Our work establishes benign overfitting beyond the nearly orthogonal data regime and can be applied to a wider class of mixtures than the prior work such as polynomially tailed mixture. Moreover, we also establish a novel lower bound of the classification error in the case of Gaussian mixture, which shows tightness of our result and helps characterize cases when benign overfitting provably fails. Even with gradient flow, benign overfitting in leaky ReLU neural networks is limited to nearly orthogonal data setting and sub-Gaussian mixture model (Frei et al. (2023b)).
@@
Most prior work on the implicit bias of (leaky) ReLU neural networks for binary classification has considered gradient flow. For gradient flow, Lyu and Li (2020) showed that any limit point of the network parameter of homogeneous neural networks trained with exponential type loss is a KKT point of a constrained optimization problem and that margin maximization occurs. Ji and Telgarsky (2020) further established directional convergence of the parameter of homogeneous neural networks trained with the exponential type loss. Taking advantage of the directional convergence of gradient flow, Lyu et al. (2021); Bui Thi Mai and Lampert (2021); Sarussi et al. (2021); Frei et al. (2023a); Min et al. (2024) obtained detailed characterizations of the convergent direction of the parameters of (leaky) ReLU two-layer neural networks under further assumptions on the training data.
In contrast to gradient flow, directional convergence of network parameters for ReLU type neural networks was not proven for gradient descent with exponential type loss, even under strong assumptions. As a result, only weaker forms of implicit bias have been obtained for gradient descent. Lyu and Li (2020) showed that margin maximization also occurs for the network parameters of a homogeneous network trained by gradient descent, but their assumptions include smoothness of the network, and hence rule out (leaky) ReLU networks. Frei et al. (2023a) showed that the stable rank of network parameters stays bounded by a constant for two-layer neural networks, but their result is limited to the smoothed approximation of leaky ReLU. Kou et al. (2023b) proved similar results for (leaky) ReLU two-layer neural networks assuming that the number of positive and negative neurons are equal. Moreover, Frei et al. (2023a); Kou et al. (2023b) are limited to the rather specialized scenario of nearly orthogonal data regime. Cai et al. (2025) showed directional convergence of gradient descent, but their result is not applicable to ReLU type neural networks since their result requires networks to be twice differentiable with respect to the parameters. Schechtman and Schreuder (2025) extended Lyu and Li (2020) by removing twice differentiability but their result does not guarantee directional convergence and assumes that training is already in the late-stage.
@@
We also note that Karhadkar et al. (2024) studied benign overfitting in leaky ReLU networks using hinge loss instead of exponential type loss.
@@
Gradient Flow Ji and Telgarsky (2020) Frei et al. (2023b) NA
@@ 3 Preliminaries
where $\boldsymbol{x}\in\mathbb{R}^{p}$, $\boldsymbol{w}_{j}\in\mathbb{R}^{p}$, $W=(\boldsymbol{w}_{1},\boldsymbol{w}_{2},\ldots,\boldsymbol{w}_{m})\in\mathbb{R}^{p\times m}$, $a_{j}=\pm\tfrac{1}{\sqrt{m}}$ and $\phi$ is $\gamma$-leaky ReLU activation, i.e. $\phi(x)=\max\{x,\gamma x\}$ for $\gamma\in(0,1)$. Letting $\zeta(x)=1$ if $x\geq 0$ and $\zeta(x)=\gamma$ if $x<0$, we may conveniently write $\phi(x)=\zeta(x)x$.
@@
For $W\in\mathbb{R}^{p\times m}$, we say benign overfitting occurs if the network $f(\cdot;W)$ classifies the training data perfectly (interpolation) while arbitrarily small test error is achieved for sufficiently large $n$.
@@ 4 Directional Convergence of Gradient Descent
We start from our main result on directional convergence of gradient descent in leaky ReLU two-layer neural networks trained on mixture data. Our result provides precise characterization of the convergent direction, which is a key for our analysis on benign overfitting. The key to prove directional convergence is neuron activation. We say that $j$-th neuron $\phi(\langle\cdot,\boldsymbol{w}_{j}\rangle)$ is activated if $a_{j}y_{i}\phi(\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle)>0$ holds for any $i\in I$, which can be equivalently written as $a_{j}y_{i}\langle\boldsymbol{x}_{i},\boldsymbol{w}_{j}\rangle>0$ for any $i\in I$.
@@
Assumption 4.4 and 4.6 allow us to carefully control $\langle y_{k}\boldsymbol{x}_{k},y_{i}\boldsymbol{x}_{i}\rangle$. Small $\varepsilon_{1},\varepsilon_{3}$ ensure that the training data are nearly orthogonal. By taking $\boldsymbol{\mu}=0$, this case includes the nearly orthogonal data regime studied in the prior work (Frei et al. (2023a); Kou et al. (2023b)). The other assumptions are analogous to Case 1.
@@
Furthermore, the convergent direction $\{\boldsymbol{\hat{w}}_{j}\}_{j\in J}$ can also be given by $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{+}$ for $j\in J_{+}$ and $\boldsymbol{\hat{w}}_{j}=\boldsymbol{w}_{-}$ for $j\in J_{-}$, which is the unique solution to the following optimization problem:
@@
We note that condition (i) goes beyond nearly orthogonal data regime studied in the existing literature since there is no upper bound to the magnitude of the signal $\boldsymbol{\mu}$, while condition (ii) contains the nearly orthogonal data regime studied in Frei et al. (2023a); Kou et al. (2023b). As discussed in Introduction, directional convergence of gradient descent in ReLU type neural networks was not established in the existing literature even under nearly orthogonal data regime. In addition, the characterizations of the convergent direction in Theorem 4.8 were obtained only for gradient flow trained on nearly orthogonal data by Frei et al. (2023a).
We also note that the optimization problem (5) has a unique solution since the objective function is strictly convex.
@@ 5 Classification Error of the Convergent Direction
where $\kappa(t)=\mathbb{P}(\xi_{1}\geq t)$ and $\beta_{min}$ is the smallest eigenvalue of $\Sigma$ and $\beta_{max}=\|\Sigma\|$.
@@ 6 Benign Overfitting on Mixture Models
    $n\geq C$, $\alpha\leq\frac{1}{8{\rm tr}(\Sigma)}$, $\sigma\leq 0.1\gamma\alpha\sqrt{\frac{{\rm tr}(\Sigma)}{m}}$, $\|\boldsymbol{\mu}\|^{2}\geq C\|\Sigma^{\frac{1}{2}}\boldsymbol{\mu}\|$,
@@
We note that $\hat{W}$ perfectly classifies the training data since the training loss $\mathcal{L}(t)$ is driven to zero by gradient descent. Thus, Theorem 6.2 implies that benign overfitting occurs if $n\|\boldsymbol{\mu}\|^{4}\gg\|\Sigma\|{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\gg\|\Sigma\|$ under (b) is satisfied. For Gaussian mixture model, we see that benign overfitting provably fails even though directional convergence occurs if $n\|\boldsymbol{\mu}\|^{4}\lesssim\beta_{min}{\rm tr}(\Sigma)$ under (a) or $\|\boldsymbol{\mu}\|^{2}\lesssim\beta_{min}$ under (b).
@@
The Bayes error rate is given by $\kappa(\|\Sigma^{-1/2}\boldsymbol{\mu}\|)$ if $\boldsymbol{z}\sim\mathcal{N}(0,\Sigma)$ and $\Sigma$ is intvertible (Section 1.2.1, Wainwright (2019)). Therefore, our bounds in the case of isotropic Gaussian mixture is tight up to a constant factor in $\kappa$ in the strong signal regime, i.e. the implicit bias of gradient descent leads to the Bayes optimal.
We also note that Karhadkar et al. (2024) studied implicit bias and benign overfitting in leaky ReLU networks in a somewhat different setting. Since their work studied hinge loss instead of exponential type loss and label-flipping noise is added, it is difficult to directly compare their result with ours. Nevertheless, their distributional setting can be viewed as a special case of ours. In fact, their distributional setting (ignoring the label-flipping noise) is essentially a Gaussian mixture with $\|\boldsymbol{\mu}\|\approx 1$ and ${\rm tr}(\Sigma)\approx 1$, which falls into the strong signal regime in our work.
@@
Consider model (PM) with $\Sigma=I_{p}$. Then, benign overfitting occurs with the convergent direction $\hat{W}$ with high probability if either one of the following holds:
@@ 7 Proof Sketch
As noted in Section 4, the key to show directional convergence of the gradient descent is activation of neurons.
@@ 8 Discussion
Acknowledgments
We thank Stanislav Volgushev and Piotr Zwiernik for helpful discussions and guidance. The author is further grateful to the anonymous reviewers and the area chair of ICLR for constructive feedback that improved the manuscript. We would also like to thank Xingyu Xu for helpful discussions regarding the work Xu and Gu (2023).
@@ Appendix B Proof of Directional Convergence
where $\kappa(t)=\mathbb{P}(\xi_{1}\geq t)$ and $\beta_{min}$ is the smallest eigenvalue of $\Sigma$ and $\beta_{max}=\|\Sigma\|$.
@@
$=\frac{1}{2}\mathbb{P}(|\langle\boldsymbol{w},\boldsymbol{z}\rangle|>\langle\boldsymbol{w},\boldsymbol{\mu}\rangle).$